A merchant wants to mix gummy worms worth $3 per pound and gummy bears worth $1.5 per pound to make 30 pounds of a mixture worth $2.10 per pound. how many pounds of each type of candy should he use.
Answer by checkley75(3666)

30*2.10=3.00X+1.50(30-X)
63=3X+45-1.5X
1.5X=63-45
1.5X=18
X=18/1.5
X=12 POUNDS OF GUMMY WORMS NEEDED
30-12=18 POUNDS OF GUMMY BEARS
PROOF
30*2.10=3*12+1.5*18
63=36+27
63=63